Can you please help me with these questions.

1). Maximize P subject to the given conditions. P = 6x − 7y x ≥ 0 y ≤ 3 x − y ≤ 4

maximum value P =

point where maximum occurs (x, y) =

2).

Find the coordinates of the vertex of the graph of the equation

y = 1/2x2 − x + 2. (x,y) =
Answer #1

Soleilion : P = 6x - Yy Maximize P. subject to constraints : 220 y<3 x-y<4 Stap 1: Couvert all the constraints into : equatioStep 3: Test of optimality Carners A(0,-4) B(7,3) P- 6x - 6x-ty P= -7-4) = 28 P = 6 XY - 7(3) = 44 -21 = 26 c(0,3) P = -7(%)